# E2CL: Exploration-based Error Correction Learning for Embodied Agents

### v1

@@ Abstract
Language models are exhibiting increasing capability in knowledge utilization and reasoning. However, when applied as agents in embodied environments, they often suffer from misalignment between their intrinsic knowledge and environmental knowledge, leading to infeasible actions. Traditional environment alignment methods, such as supervised learning on expert trajectories and reinforcement learning, face limitations in covering environmental knowledge and achieving efficient convergence, respectively. Inspired by human learning, we propose Exploration-based Error Correction Learning (E2CL), a novel framework that leverages exploration-induced errors and environmental feedback to enhance environment alignment for LM-based agents. E2CL incorporates teacher-guided and teacher-free exploration to gather environmental feedback and correct erroneous actions. The agent learns to provide feedback and self-correct, thereby enhancing its adaptability to target environments. Evaluations in the Virtualhome environment demonstrate that E2CL-trained agents outperform those trained by baseline methods and exhibit superior self-correction capabilities.
@@
Hanlin Wang, Chak Tou Leong, Jian Wang, Wenjie Li Department of Computing, The Hong Kong Polytechnic University {hanlin-henry.wang,chak-tou.leong,jian-dylan.wang}@connect.polyu.hk cswjli@comp.polyu.edu.hk
@@ 1 Introduction
Language Models (LMs) are becoming increasingly capable of knowledge utilization and reasoning across various knowledge-intensive tasks (Yao et al., 2022; Lewkowycz et al., 2022; Hao et al., 2023). This success motivates researchers to apply LMs to build LM-based agents in embodied environments, which similarly requires the use of reasoning and planning upon environmental knowledge (Li et al., 2022; Xiang et al., 2024). In this case, LM-based agents are asked to plan appropriate actions based on the given environmental information and the history of actions already taken. However, the knowledge acquired by these LM-based agents comes from general-purpose corpora during pre-training, and as a result the intrinsic knowledge of these models often misalign with environmental knowledge. Such environmental knowledge involves physical constraints that LMs have not yet explored. For example, if the embodied agent holds two objects, it is prohibited to grab one more other object. This misalignment causes LM-based agents to frequently infer actions that cannot be executed in the environment, hindering their application in real-world environments.
@@
Figure 1: Traditional "Success Only Learning" relies on imitating provided expert behaviors, limiting comprehensiveness. Our proposed framework, Exploration-based Error Correction Learning (E2CL), enhances learning by incorporating exploration-induced errors and environmental feedback during training, leading to better alignment with target environments. During inference, the agent utilizes the learned ability to conduct self-feedback for continuous self-correction.
@@
To address this issue, two primary types of environment alignment methods have been explored. The first type involves having LM-based agents undergo supervised learning on expert trajectories, which are human-labeled sequences of observations and actions (Li et al., 2022; Chen et al., 2023). Nevertheless, these trajectories often fail to fully cover the knowledge within the environment, such as scenarios where certain actions cannot be executed. The second type is based on reinforcement learning, which allows agents to freely explore the environment, collect trajectories that comprehensively cover the environment’s knowledge, and obtain rewards based on these trajectories’ success or failure (Tan et al., 2024; Carta et al., 2023). However, the rewards are sparsely obtained because the performance evaluation of the agent is based on a complete trajectory. This makes the learning process difficult to converge.
@@
Human learning is not comprehensive or efficient if it relies solely on imitating experts’ behavior or merely knowing whether an action is correct. Instead, by collecting and understanding feedback from the environment via exploration and learning to correct errors based on the feedback, humans can learn comprehensively and efficiently. Inspired by this, we propose a novel exploration framework for LM-based agents to align with environments, which is called Exploration-based Error Correction Learning (E2CL). As depicted in Figure 1, our framework incorporates exploration-induced errors and environmental feedback, leading to a comprehensive alignment with target environments.
@@
During the process of E2CL, we ask a pretrained agent to perform predefined tasks and explore the environment to collect experiences in both efficient and comprehensive manners. This is achieved by two different proposed schemes, namely teacher-guided exploration and teacher-free exploration. The former prompts the agent to perform one-step exploration given sliced expert trajectories, whereas the latter allows the agent to continue exploring until it infers a stop. In these two exploration phases, we collect the feedback given by the environment when the agent makes errors, as well as the correct actions corresponding to these error actions. Having these exploration trajectories with additional correction, we train the agent to provide feedback on their trajectories and correct their error actions based on the feedback. To apply learned self-correction ability, we further propose Speculative Inference, which performs corrections if the initial planned actions are inferred to be errors by the agent’s feedback.
@@
In summary, our main contributions are as follows. (1) We introduce the Exploration-based Error Correction Learning (E2CL) framework, enabling LM-based agents to align with environments through effective feedback-driven exploration and correction. (2) We propose two novel exploration schemes, teacher-guided and teacher-free exploration, that facilitate the collection of correction and feedback via agent-environment interaction. (3) We introduce a novel action inference algorithm, namely speculative inference, which can prevent executable errors from occurring. (4) We demonstrate the superior performance of E2CL-trained agents in the Virtualhome environment, surpassing baseline methods and showcasing the potential of our approach for real-world deployment.
@@ 2 Method
In this section, we introduce our framework, E2CL. This framework equips LM-based agents with self-feedback and self-correction capabilities, which enhances their ability to tackle tasks in new environments.
@@ 2.1 Task Formulation
Figure 2: The illustration of the proposed Exploration-based Error Correction Learning (E2CL) framework.
@@ 2.2 Exploration-based Error Correction Learning
Our E2CL framework consists of three phases of learning and exploration within the environment: the pre-tuning phase, the exploration phase, and the training phase. In the pre-tuning phase, we equip the agent with basic planning ability before exploration. Then, in the exploration phase, the agent collects exploration experience in the environment via two complementary schemes, as shown in Figure 2. Following this, in the training phase, the agent is trained to align with the environmental knowledge from the collected experience and to develop the ability to provide feedback and correct its own errors.
@@
To serve as the foundation for environmental exploration, we aim to empower LM-based embodied agents with a basic planning capability. Given a dataset $\mathcal{J}=\{(q^{i}_{p},j^{i}_{n_{i}})\}^{|\mathcal{J}|}$ with $|\mathcal{J}|$ task instructions and expert trajectories where each trajectory has $n_{i}$ steps, we first construct a planning dataset $D_{p}$ by slicing each trajectory into sub-trajectories of varying lengths from $1$ to $n_{i}$. Formally, the planning dataset $D_{p}$ is defined as: $D_{p}=\bigcup_{i=1}^{|\mathcal{J}|}\bigcup_{t=1}^{n_{i}}\left\{(q^{i}_{p},j^{i}_{t})\mid j^{i}_{t}\subseteq j^{i}_{n_{i}}\land(q^{i}_{p},j^{i}_{n_{i}})\in\mathcal{J}\right\}$. Notably, we sample a subset of $D_{p}$, denoted as $D_{p^{\prime}}$, for pre-tuning to avoid overfitting to expert trajectories and maintain exploration diversity. Then, we fine-tune the LM-based agent by minimizing neg-likelihood loss:
@@
Notably, $a_{t}$ consists of multiple tokens. Therefore, when calculating the loss, it effectively becomes an auto-regressive loss over a sequence of tokens, following previous practices. This approach is consistently applied in the latter stages of training as well.
@@
During TFE, the agent iterates through each task instruction $q_{p}\sim\mathcal{Q}$ and accrodingly obtain trajectories $j_{t}=(\hat{a}_{1},\hat{o}_{1},\dots,\hat{a}_{t},\hat{o}_{t})$. Similar to TGE, whenever the agent predicts a non-executable $\hat{a}_{t}$, the environment provides feedback $f_{t}$ indicating why this action is non-executable. To obtain the executable action at this step without manual intervention, we leverage an LLM with powerful reasoning ability (e.g., GPT-4o) to automatically correct the action, yielding $a_{t}$. Considering the LLM may not always provide perfect corrections due to a lack of environment alignment, we further filter its predictions to ensure the corrections are executable. As a result, we obtain the feedback dataset $D_{f}^{\text{TFE}}$ and $D_{c}^{\text{TFE}}$, each in the same form as the samples in $D_{f}^{\text{TGE}}$ and $D_{c}^{\text{TGE}}$, respectively.
Training Phase
After the above two phases, we obtain the planning dataset $D_{p}$, the feedback dataset $D_{f}=D_{f}^{\text{TGE}}\bigcup D_{f}^{\text{TFE}}$, and the correction dataset $D_{c}=D_{c}^{\text{TGE}}\bigcup D_{c}^{\text{TFE}}$.
Next, we train the agent to align with the environmental knowledge gathered from these datasets and to develop the ability to provide feedback and correct its own errors. This is achieved by fine-tuning the agent to minimize the following losses:
We refer the reader to the pseudo-code of the overall E2CL process in Appendix F.
2.3 Speculative Inference
To utilize the learned abilities in the training phase, we propose speculative inference, a process of predicting the error occurring ahead and correcting itself. This process reduces execution errors and generates correct action based on self-generated feedback.
To be more precise, when given each test task instruction $q_{p}$, the agent initially predicts a action $\hat{a}_{t}\sim\pi_{\theta}(\cdot|q_{p},j_{t-1})$. However, this action $\hat{a}_{t}$ will not be executed immediately. The agent will ‘reflect’ itself and generate an environment feedback $\hat{f}_{t}\sim\pi_{\theta}(\cdot|q_{f},j_{t-1},\hat{a}_{t})$. If the agent believes the initial action $\hat{a}_{t}$ is executable, then this action will be executed. Otherwise, the embodied agent will correct this action $\hat{a}$ and predict a new action $\hat{a}_{c}\sim\pi_{\theta}(\cdot|q_{c},j_{t-1},\hat{a}_{t},\hat{f}_{t})$. Once the corrected action $\hat{a}_{c}$ passes its own check, this action will be executed at this step and concatenated into trajectory $\hat{j}_{t-1}$. The above process is iterated until the agent assumes the task is finished or the total steps exceed the maximum threshold. The process of speculative inference for each test task is shown in  Algorithm 1.
Input: $\pi_{\theta}$: The policy of robot agent, VS: Virtualhome Simulator
@@ 2.3 Speculative Inference
return R: Execution result for each task
@@ 3.1 Experimental Settings
VirtualHome Environment & Tasks
We evaluate our method in the VirtualHome Environment (Puig et al., 2018), a simulation platform for performing typical household activities.
During the experiment, we are able to access data, such as states of objects, observation of agents, and environmental feedback, from the simulator. We refer to Appendix D for more details of the environment.
We use the predefined tasks from ActivityPrograms Knowledge Base (Puig et al., 2018) for the experiment. It contains 292 unique high-level household tasks, with 1374 unique action plans and 6201 unique environmental settings in total extracted from VirtualHome.
After filtering low-quality tasks, we conduct experiments on a total of 285 tasks. They are randomly divided into a training set of 235 tasks and a test set of 50 tasks. We select 50 tasks from the training set as seen tasks, while the 50 tasks in the test set as unseen tasks. We evaluate the method on both seen tasks and unseen tasks.
@@
We compare our method with both prompting-based methods and other tuning-based baseline methods. Similar to our approach, tuning-based methods achieve alignment between the embodied agent and the environment via model fine-tuning. (1) Language-planner (Huang et al., 2022) aims to inject environment knowledge into prompt and prompted Large Language Models to output action. We utilize GPT-4o as the foundational model for our baseline method. (2) We perform Behavior Cloning (BC) on expert planning data (Chen et al., 2023; Zeng et al., 2023), which is the same method used in the pre-tuning phase of our methods and other baselines. (3) LWM (Xiang et al., 2024) employs a robot agent to interact with the environment and collect a large amount of environmental knowledge data to fine-tune the model. (4) Plasma (Brahman et al., 2023) leverages ChatGPT to generate multi-task planning-related data for model training. (5) Lema (An et al., 2023) enhances the agent’s reasoning capabilities by providing error-correction data pairs during model fine-tuning. (6) NAT (Wang et al., 2024b) implements a negative-aware training approach, enabling LM-based agents to effectively learn from both positive and negative examples.
@@
Following previous studies (Puig et al., 2018; Raman et al., 2022), we evaluate our action plans across three metrics: executability (Exec.), affordance rate (AR), and longest common sequence (LCS). Executability measures whether an action plan can be correctly parsed and satisfies the common-sense constraints of the environment. Specifically, the parsed action must contain only allowable action, and the objects must be in the environment. Moreover, the action must satisfy the pre-conditions (e.g., the embodied agent cannot send email before walking to the computer) and post-conditions (e.g., the state of TV changes from closed to open after the agent opens it). Similarly to executablility, affordance rate measures the average percentage of all plan steps that are executable, in cases where the entire plan is not executable. However, executability and affordance rate only can reflect whether the agent could compliant with environment physical constraints, but they cannot reflect whether the plan is correct. LCS calculates the length of the longest common subsequence between generated plans and the ground truth plans, normalized by the maximum length of the two.
3.2 Results
As shown in  Table 1, our method outperforms both prompting-based and tuning-based baseline methods across multiple metrics, showing the superiority of our method.
Notably, the prompting-based method significantly lags behind all tuning-based methods in different metrics. Despite this contradicts with the experience that LLMs exhibit exceptional general reasoning capabilities, we observe that the actions generated by prompt-based methods, while seemingly reasonable, fail to comply with the physical constraints of the environment often. Regarding tuning-based baseline methods, our method demonstrates significant improvements over BC in both seen and unseen tasks.
Moreover, LWM and Plasma, which are also fed by expert planning data and can be seen as augmented versions of BC, only show a marginal increase in performance. Compared to these BC-based methods, the method utilizing failure data, i.e., Lema and NAT, demonstrates better performance. Taking a step further, we evolve this idea by training the agent to develop self-feedback and self-correction capabilities through its failure experiences. The results show that our method increases executability-related metrics by up to 15% and LCS by up to 10% compared with Lema and NAT. This demonstrates that these two capabilities effectively enable the agent to align with the environment for task-solving.

### v2

@@ Abstract
Language models are exhibiting increasing capability in knowledge utilization and reasoning. However, when applied as agents in embodied environments, they often suffer from misalignment between their intrinsic knowledge and environmental knowledge, leading to infeasible actions. Traditional environment alignment methods, such as supervised learning on expert trajectories and reinforcement learning, encounter limitations in covering environmental knowledge and achieving efficient convergence, respectively. Inspired by human learning, we propose Exploration-based Error Correction Learning (E2CL), a novel framework that leverages exploration-induced errors and environmental feedback to enhance environment alignment for embodied agents. E2CL incorporates teacher-guided and teacher-free explorations to gather environmental feedback and correct erroneous actions. The agent learns to provide feedback and self-correct, thereby enhancing its adaptability to target environments. Extensive experiments in the VirtualHome environment demonstrate that E2CL-trained agents outperform those trained by baseline methods and exhibit superior self-correction capabilities.
@@
Hanlin Wang*,   Chak Tou Leong*,   Jian Wang,   Wenjie Li Department of Computing, The Hong Kong Polytechnic University {hanlin-henry.wang,chak-tou.leong,jian-dylan.wang}@connect.polyu.hk cswjli@comp.polyu.edu.hk
footnotetext: *Equal contribution.
@@ 1 Introduction
Language Models (LMs) are becoming increasingly capable of knowledge utilization and reasoning across various knowledge-intensive tasks (Yao et al., 2022; Lewkowycz et al., 2022; Hao et al., 2023). This success motivates researchers to build LM-based agents in embodied environments, which similarly requires the use of reasoning and planning upon environmental knowledge (Li et al., 2022; Xiang et al., 2024). In this case, LM-based agents are asked to plan appropriate actions based on the given environmental information and the history of actions already taken. However, the knowledge acquired by these LM-based agents comes from general-purpose corpora during pre-training, and as a result the intrinsic knowledge of these models often misalign with environmental knowledge. Such environmental knowledge involves physical constraints that LMs have not yet explored. For example, if the embodied agent holds two objects, it is prohibited to grab one more other object. This misalignment causes LM-based agents to frequently produce actions that cannot be executed in the environment, hindering their applications in real-world environments.
@@
Figure 1: Traditional “success only learning” relies on imitating provided expert behaviors, limiting comprehensiveness. Our proposed exploration-based error correction learning (E2CL) framework enhances learning by incorporating exploration-induced errors and environmental feedback during training, leading to better alignment with target environments. During inference, the agent utilizes the learned abilities to conduct self-feedback for continuous self-correction.
@@
To address the above issue, two primary types of environment alignment methods have been explored. The first type involves having LM-based agents undergo supervised learning on expert trajectories (Li et al., 2022; Chen et al., 2023), which are human-labeled sequences of observations and actions. Nevertheless, these trajectories often fail to fully cover the knowledge within the environment, such as scenarios where certain actions cannot be executed. The second type is based on reinforcement learning (RL), which allows agents to freely explore the environment, collect trajectories that comprehensively cover the environment’s knowledge, and obtain rewards based on these trajectories’ success or failure (Tan et al., 2024; Carta et al., 2023). However, the rewards are sparsely obtained because the performance evaluation of the agent is based on a complete trajectory. This makes the learning process difficult to converge.
@@
Human learning is not comprehensive nor efficient if it relies solely on imitating experts’ behavior or merely knowing whether an action is correct. Instead, by collecting and understanding feedback from the environment via exploration and learning to correct errors based on the feedback, humans can learn comprehensively and efficiently. Inspired by this, we propose a novel exploration framework for LM-based agents to align with environments, which is called Exploration-based Error Correction Learning (E2CL). As depicted in Figure 1, our framework incorporates exploration-induced errors and environmental feedback, leading to a comprehensive alignment with target environments.
@@
In detail, we adopt a pretrained model as the agent to perform predefined tasks and explore the environment to collect experiences in both efficient and comprehensive manners. This is achieved by two different proposed schemes, namely teacher-guided exploration and teacher-free exploration. The former prompts the agent to perform one-step exploration given sliced expert trajectories, whereas the latter allows the agent to continue exploring until it infers a stop. In these two exploration phases, we collect the feedback given by the environment when the agent makes errors, as well as the correct actions corresponding to these error actions. Having these exploration trajectories with additional correction, we train the agent to provide feedback on their trajectories and correct their error actions based on the feedback. To utilize the learned self-correction ability, we further propose a speculative inference algorithm, which performs corrections if the initially planned actions are inferred to be errors according to the feedback from the agent.
@@
In summary, our main contributions are as follows. (1) We introduce the Exploration-based Error Correction Learning (E2CL) framework, enabling LM-based agents to align with environments through effective feedback-driven exploration and correction. (2) We propose two novel exploration schemes, teacher-guided and teacher-free explorations, which facilitate the collection of correction and feedback via agent-environment interactions. (3) We introduce a novel action inference algorithm called speculative inference, which effectively avoids executable errors. (4) We demonstrate the superior performance of E2CL-trained agents in the VirtualHome environment, surpassing baseline methods and showcasing the potential of our approach for real-world deployment.
@@ 2 Method
In this section, we propose an exploration-based error correction learning (E2CL) framework. This framework focuses on equipping LM-based agents with self-feedback and self-correction capabilities. The overview of E2CL is depicted in Figure 2.
@@ 2.1 Task Formulation
Figure 2: Overview of the proposed Exploration-based Error Correction Learning (E2CL) framework.
@@ 2.2 Exploration-based Error Correction Learning
Our E2CL framework consists of three phases of learning and exploration within the environment: pre-tuning, exploration, and training. In the pre-tuning phase, the agent is equipped with basic planning ability before exploration. Then, the agent collects exploration experiences in the environment via two complementary schemes, as shown in Figure 2. Following this, in the training phase, the agent is trained to align with the environmental knowledge from the collected experiences. With this alignment, the agent is expected to provide feedback by itself and correct its own errors.
@@
To serve as the foundation for environmental exploration, we aim to empower LM-based embodied agents with basic planning capabilities. Given a dataset $\mathcal{J}=\{(q^{i}_{p},j^{i}_{n_{i}})\}^{|\mathcal{J}|}$ with $|\mathcal{J}|$ task instructions and expert trajectories where each trajectory has $n_{i}$ steps, we first construct a planning dataset $D_{p}$ by slicing each trajectory into sub-trajectories of varying lengths from $1$ to $n_{i}$. Formally, the planning dataset $D_{p}$ is defined as:
$D_{p}=\bigcup_{i=1}^{|\mathcal{J}|}\bigcup_{t=1}^{n_{i}}\left\{(q^{i}_{p},j^{i}_{t})\mid j^{i}_{t}\subseteq j^{i}_{n_{i}}\land(q^{i}_{p},j^{i}_{n_{i}})\in\mathcal{J}\right\}$ (1)
Notably, we sample a subset of $D_{p}$, denoted as $D_{p^{\prime}}$, for pre-tuning to avoid overfitting to expert trajectories and maintain exploration diversity. Then, we fine-tune the LM-based agent by minimizing negative log-likelihood loss:
@@
where $a_{t}$ denotes an action consisting of multiple tokens. Therefore, when calculating the loss, it effectively becomes an auto-regressive loss over a sequence of tokens, following previous practices. This approach is consistently applied in the latter stages of training as well.
@@
During TFE, the agent iterates through each task instruction $q_{p}\sim\mathcal{Q}$ and accrodingly obtain trajectories $j_{t}=(\hat{a}_{1},\hat{o}_{1},\dots,\hat{a}_{t},\hat{o}_{t})$. Similar to TGE, whenever the agent predicts a non-executable $\hat{a}_{t}$, the environment provides feedback $f_{t}$ indicating why this action is non-executable. To obtain the executable action at this step without manual intervention, we leverage an LLM with powerful reasoning ability (e.g., GPT-4o) to automatically correct the action, yielding $a_{t}$. Considering the LLM may not always provide perfect corrections due to a lack of environment alignment, we further filter its predictions to ensure the corrections are executable. Specifically, the corrected action $a_{t}$ predicted by LLM would replace original error action $\hat{a}_{t}$ and be executed in the environment. If the corrected action $a_{t}$ is successful executable in the environment, we would collect these correction data into $D_{c}^{\text{TFE}}$. As a result, we obtain the feedback dataset $D_{f}^{\text{TFE}}$ and $D_{c}^{\text{TFE}}$, each in the same form as the samples in $D_{f}^{\text{TGE}}$ and $D_{c}^{\text{TGE}}$, respectively.
Input: $\pi_{\theta}$: policy of the embodied agent, ES: environment simulator
@@
return R
@@
Training Phase
After the above two phases, we obtain the planning dataset $D_{p}$, the feedback dataset $D_{f}=D_{f}^{\text{TGE}}\bigcup D_{f}^{\text{TFE}}$, and the correction dataset $D_{c}=D_{c}^{\text{TGE}}\bigcup D_{c}^{\text{TFE}}$. Next, we train the agent to align with the environmental knowledge gathered from these datasets and to develop the ability to provide feedback and correct its own errors. This is achieved by fine-tuning the agent to minimize the following losses:
We refer the reader to the pseudo-code of the overall E2CL process in Appendix E.
2.3 Speculative Inference
To utilize the learned abilities in the training phase, we propose speculative inference algorithm, a process of inferring errors that may occur ahead of execution and correcting the errors by itself. Based on self-produced feedback, the agent is desired to reduce execution errors and generate correct actions.
To be more precise, when given each test task instruction $q_{p}$, the agent initially predicts an action $\hat{a}_{t}\sim\pi_{\theta}(\cdot|q_{p},j_{t-1})$. However, this action $\hat{a}_{t}$ will not be executed immediately. The agent will “reflect” itself and generate an environment feedback $\hat{f}_{t}\sim\pi_{\theta}(\cdot|q_{f},j_{t-1},\hat{a}_{t})$. If the agent believes the initial action $\hat{a}_{t}$ is executable, then this action will be executed. Otherwise, the agent will correct this action $\hat{a}$ and predict a new action $\hat{a}_{c}\sim\pi_{\theta}(\cdot|q_{c},j_{t-1},\hat{a}_{t},\hat{f}_{t})$. Once the corrected action $\hat{a}_{c}$ passes its own check, this action will be executed at this step and concatenated into trajectory $j_{t-1}$. The above process is iterated until the agent assumes the task is finished or the total steps exceed the maximum threshold. The process of speculative inference for each test task is shown in  Algorithm 1.
@@ 3.1 Experimental Settings
Embodied Environment & Tasks
In this work, we aim to bridge the misalignment gap between LM-based agents and environmental physical constraints, and our method focuses on learning to correct erroneous actions. This implies that providing environmental feedback regarding detailed error information is a necessity. To the best of our knowledge, VirtualHome (Puig et al., 2018), which focuses on performing typical household tasks, is the most suitable simulated environment for exploration. In VirtualHome, the environmental feedback is in the form of a textual message regarding the physical environmental constraints. If an agent’s action is executable, the environment will return a short message of “True”, otherwise, the environment will return an error message indicating why the action is not executable. For example, one of the typical error types for such feedback is called “Missing Object”, which arises when the agent is not holding the necessary object to complete an action. More error types of environmental feedback are illustrated in Appendix C.1.
We use the predefined tasks from ActivityPrograms  (Puig et al., 2018) knowledge base for the experiment. It contains 292 unique high-level household tasks, with 1374 unique action plans and 6201 unique environmental settings in total extracted from VirtualHome. After filtering low-quality tasks, we conduct experiments on a total of 285 tasks. They are randomly divided into a training set of 235 tasks and a test set of 50 tasks. We select 50 tasks from the training set as seen tasks, while the 50 tasks in the test set as unseen tasks. We evaluate the method on both seen tasks and unseen tasks.
Note that during the experiment, we are able to access environmental feedback from the environment simulator. In the testing process, to align with real-world conditions, we do not expect the agents to access such environmental feedback. We refer to Appendix B for more details of the environment simulator.
@@
We compare our method with both prompting-based methods and other tuning-based baseline methods. Similar to our approach, tuning-based methods achieve alignment between the embodied agent and the environment via model fine-tuning. (1) Language-planner (Huang et al., 2022) aims to inject environment knowledge into prompt and prompted Large Language Models to output action. To better represent prompting-based methods, we use the most powerful LLM, GPT-4o, as the foundation model for our baseline method. (2) We perform Behavior Cloning (BC) on expert planning data (Chen et al., 2023; Zeng et al., 2023), which is the same method used in the pre-tuning phase of our methods and other baselines. (3) We conduct Proximal Policy Optimization (PPO) (Schulman et al., 2017) after BC. Similar to VirtualHome (Puig et al., 2018), we utilize LCS as the reward for RL training. (4) LWM (Xiang et al., 2024) employs an embodied agent to interact with the environment and collect a large amount of environmental knowledge data to fine-tune the model. (5) Plasma (Brahman et al., 2023) leverages ChatGPT to generate multi-task planning-related data for model training. (6) Lema (An et al., 2023) enhances the agent’s reasoning capabilities by providing error-correction data pairs during model fine-tuning. (7) NAT (Wang et al., 2024b) implements a negative-aware training approach, enabling LM-based agents to effectively learn from both positive and negative examples.
@@
Following previous studies (Puig et al., 2018; Raman et al., 2022), we evaluate our action plans across three metrics: executability (Exec.), affordance rate (AR), and longest common sequence (LCS). Executability measures whether an action plan can be correctly parsed and satisfies the common-sense constraints of the environment. Specifically, the parsed action must contain only allowable action, and the objects must be in the environment. Moreover, the action must satisfy the pre-conditions (e.g., the embodied agent cannot send email before walking to the computer) and post-conditions (e.g., the state of TV changes from closed to open after the agent opens it). Similarly to executablility, affordance rate measures the average percentage of all plan steps that are executable, in cases where the entire plan is not executable. However, executability and affordance rate only can reflect whether the agent could compliant with environment physical constraints, but they cannot reflect whether the plan is correct. LCS calculates the length of the longest common subsequence between generated plans and the ground truth plans, normalized by the maximum length of the two. It can reflect the correctness of the plans generated by the agents.
3.2 Experimental Results
Our experimental results are shown in  Table 1. As can be seen, the prompting-based method significantly lags behind all tuning-based methods in different metrics. Despite this contradicts with the experience that LLMs exhibit exceptional general reasoning capabilities, we observe that the actions generated by prompt-based methods, while seemingly reasonable, fail to comply with the physical constraints of the environment often.
Regarding tuning-based baseline methods, our E2CL method demonstrates significant improvements over BC in both seen and unseen tasks. Even with PPO applied on top of BC, the performance remains weak. This is likely because the action space is too large and the rewards are sparse, making it difficult to optimize the model in such an embodied environment. Moreover, LWM and Plasma, which are also fed by expert planning data and can be seen as augmented versions of BC, only show a marginal increase in performance. Compared to these BC-based methods, the method utilizing failure data, i.e., Lema and NAT, demonstrates better performance. Taking a step further, we evolve this idea by training the agent to develop self-feedback and self-correction capabilities through its failure experiences. The results show that our method increases executability-related metrics by up to 15% and LCS by up to 10% compared with Lema and NAT. This demonstrates that these two capabilities effectively enable the agent to align with the environment for task-solving.